• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 08:08:59 on September16,2005

    Entry number 154859

    keyword=Owl Shift Summary

    Owl Shift Summary
    ----------------------
    
    Summary: Controlled access for right Q1 for the first half of the shift, and no 
    beam due to the Hall A laser being unlocked.for the last half of the shift.  
    Right Q1 will continue to be worked on during day shift.
    
    -----------------
    
    00:30 - Controlled Access to fix right Q1.  Riad and Gary went down to the hall.
    
    01:30 - Mark Stevens is coming in to help Gary with Q1.
    
    04:00 - Q1 can't be fixed right now, so we'll try to do the profile scan test.
    
    04:15 - MCC can't get the Hall A laser to lock.
    
    04:30 - Hall A laser won't be worked on until Day shift, so we take pedestal 
    data for the rest of the morning
    
    06:35 - Asked for Controlled Access since MCC can't deliver beam, and the techs 
    are coming in at 08:00 to work on Q1.
    
    SUMMARY of RUNS
    -----------------
    5080 - pedestal run
    5081 - pedestal run
    5082 - pedestal run
    5083 - pedestal run
    5084 - pedestal run
    5085 - pedestal run
    5086 - pedestal run
    5087 - pedestal run